OpenStudy (anonymous):

Write the equation of a hyperbola with a center at (-5, -3), vertices at (-5, -5) and (-5, -1) and co-vertices at (-11, -3) and (1, -3).

OpenStudy (amistre64):

id move it to the origin, and then adjust it back to center

OpenStudy (amistre64):

Write the equation of a hyperbola with a center at (-5+5, -3+3), vertices at (-5+5, -5+3) and (-5+5, -1+3) co-vertices at (-11+5, -3+3) and (1+5, -3+3).

OpenStudy (reemii):

it is a vertical hyperbola, because the vertices are above and below the center. Equation will be: \(\frac{(y-c_1)^2}{a^2} - \frac{(x-c_1)^2}{b^2} = 1\). The center of the parabola is \((c_1,c_2)\). \(a\)=distance between center and vertex. \(b\)=distance between center and covertex. Replace in the formula.

OpenStudy (reemii):

hyperbola*

OpenStudy (anonymous):

Thanks!

OpenStudy (reemii):

i mean \((y-c_2)^2\) *

OpenStudy (reemii):

write here what you find?

OpenStudy (anonymous):

Would it be (y +5)^2 / 4 for the first part?

OpenStudy (reemii):

almost, it's (y+3)^2 /4, you must the y-coordinate in the (y-..)^2 parenthesis.

OpenStudy (reemii):

it is then clear that it will be (x+5)^2/36 for the other part.

OpenStudy (anonymous):

Okay, thanks (:

OpenStudy (reemii):

yw